OpenStudy (anonymous):

please help me with factoring 216+m^3

OpenStudy (anonymous):

First, let's arrange it onto the way we like to see it. Due to the communitive property of addition 216+m^3 is the same thing as m^3 +216

OpenStudy (anonymous):

This is a sum of cubes, where you can take the 3rd root of the first term and the 2nd term. This is actually an equation where we use a formula \[A ^{3}+B ^{3}\rightarrow \left( A +B \right)\left( A ^{2}-AB +B ^{2} \right)\]

OpenStudy (anonymous):

The cube root of M^3 is M, the cube root of 216 is 6 So, A=m B=6 \[\left( M +6 \right)\left( M ^{2}-6m +36 \right)\]

OpenStudy (anonymous):

That's your final answer.

OpenStudy (anonymous):

um ok thank you